### v2

# Unified Stochastic Framework for Neural Network Quantization and Pruning

Abstract.
Quantization and pruning are two essential techniques for compressing neural networks, yet they are often treated independently, with limited theoretical analysis connecting them. This paper introduces a unified framework for post-training quantization and pruning using stochastic path-following algorithms. Our approach builds on the Stochastic Path Following Quantization (SPFQ) method, extending its applicability to pruning and low-bit quantization, including challenging 1-bit regimes. By incorporating a scaling parameter and generalizing the stochastic operator, the proposed method achieves robust error correction and yields rigorous theoretical error bounds for both quantization and pruning as well as their combination.
1 Department of Mathematics, University of California-San Diego, La Jolla, California 92093, USA. (haz053@ucsd.edu)
2 Department of Mathematics, Halıcıoğlu Data Science Institute, University of California-San Diego, La Jolla, California 92093, USA. (rsaab@ucsd.edu)
$\dagger$ Corresponding author
1. Introduction
Modern deep neural networks (DNNs) have achieved significant success but require substantial memory and computation due to their large number of parameters. Model compression techniques, including quantization, pruning, knowledge distillation, and low-rank decomposition, mitigate these issues. We focus on quantization and pruning.
Quantization replaces the weights in a DNN with elements from a finite set, allowing them to be represented with fewer bits and consequently not only compresses the network but also accelerates inference. DNN quantization methods include quantization-aware training and post-training quantization. Quantization-aware training (e.g., [3, 4, 7, 14, 26, 28, 33]) modifies standard neural network training to account for weight quantization, often requiring retraining and hyperparameter tuning, which can be computationally expensive, especially if one needs to repeat the process for multiple bit depths. In contrast, post-training quantization (e.g., [5, 6, 10, 13, 18, 20, 21, 27, 30, 32]) operates on pre-trained models, using relatively little data to replace floating-point parameters with quantized counterparts. Post-training quantization methods are generally much less demanding than the original training process, allowing for easy re-quantization of a base model at various bit depths.
Meanwhile, pruning techniques aim to set as many of a DNN’s weights to zero as possible. Pruning techniques can also be categorized into those that operate pre-training, during training, or post-training. Pre-training pruning methods (e.g., [15, 23, 24, 25]) prune the model before training and often rely significantly on good initialization. Methods that prune during training (e.g., [8, 12, 11, 17, 31]) attempt to identify and remove unnecessary parameters during training, and can increase overall training cost. Post-training pruning methods (e.g., [2, 9, 16, 19]) prune already trained models and fine-tune them if necessary. Similar to their quantization counterparts, post-training methods allow models to be pruned to various sparsity levels, enabling implementation on platforms with different memory, power, and computational resources.
In this theory-focused paper, we extend post-training algorithms for quantization and pruning by building on the work presented in [29], which introduced Stochastic Path Following Quantization (SPFQ). SPFQ is a stochastic error-correcting algorithm that sequentially quantizes a neuron’s weights to minimize accumulated error, with an accompanying theoretical upper bound on the error that scales as $\sqrt{\log(N)}$, where $N$ is the dimension of the neuron. This scaling compares favorably to the $\sqrt{N}$ scaling associated with other techniques. However, SPFQ’s focus on minimizing accumulated error without constraints limits its applicability to extremely low-bit quantizers, as it cannot guarantee rigorous error bounds in such cases. Additionally, while SPFQ includes an analysis for finite-bit quantization when the number of bits is not too small, its framework does not readily extend to pruning.
We provide some important generalizations to SPFQ and its analysis. We show that one can replace the quantization operator in SPFQ by any stochastic operator that satisfies certain properties, and we introduce a scaling parameter into the argument of this operator. Loosely speaking, this scaling parameter controls the amount of error correction in the algorithm, and is critical to allowing us to handle the low-bit (even 1-bit) quantization regime. These modifications make our approach versatile for various tasks, including pruning and the afore-mentioned quantization, and we prove rigorous error bounds for these applications.
This paper is organized as follows. Section 2 introduces notation. In Section 3, we present our algorithm and the intuition behind it. Section 4 states and proves the main theoretical result. Sections 5, 6, and 7 discuss applications —quantization, pruning, and combined quantization with pruning—and apply the main theorem to derive error bounds, demonstrating the algorithm’s versatility for low-bit quantization and sparsification.
2. Notation
We define an $L$-layer perceptron $\Phi:\mathbb{R}^{N_{0}}\rightarrow\mathbb{R}^{N_{L}}$ via its action
$\Phi(x)=\rho^{(L)}\circ A^{(L)}\circ\rho^{(L-1)}\circ A^{(L-1)}\circ\dots\circ\rho^{(1)}\circ A^{(1)}(x),$
[…]
We denote by $\Phi^{(i)}$ the composition of the first $i$ layers of the model, so that
$\Phi^{(i)}(x)=\rho^{(i)}\circ A^{(i)}\circ\rho^{(i-1)}\circ A^{(i-1)}\circ...\circ\rho^{(1)}\circ A^{(1)}(x)$
represents the activations at layer $i$. Similarly, we denote the compressed model by $\widetilde{\Phi}$, and its first $i$ layers by $\widetilde{\Phi}^{(i)}$. $X\in\mathbb{R}^{m\times N_{0}}$ represents the data used for compressing the network, where each row is a data point in $\mathbb{R}^{N_{0}}$. We extend notation so $\Phi^{(i)}(X)\in\mathbb{R}^{m\times N_{i}}$ denotes the application of $\Phi^{(i)}$ to each row of $X$.
For a matrix $M\in\mathbb{R}^{m\times n}$, $M_{j}\in\mathbb{R}^{m}$ represents the $j$-th column of $M$, $M_{ij}$ represents the $(i,j)$-th entry of $M$, and for a vector $v\in\mathbb{R}^{n}$, $v_{j}\in\mathbb{R}$ represents the $j$-th entry of $v$. If $A$ and $B$ are two symmetric matrices, $A\preceq B$ means $B-A$ is positive semi-definite. For vectors $x\in\mathbb{R}^{d}$, we require the $\ell_{2}$ and $\ell_{\infty}$ norms $\|x\|_{\infty}:=\max_{i}|x_{i}|$, $\|x\|:=\sqrt{\sum_{i}x_{i}^{2}}$. Meanwhile for matrices $X$, $\|X\|_{F}=\sqrt{\sum_{i,j}X_{ij}^{2}}$ denotes its Frobenius norm. We use $\mathcal{N}(\mu,\Sigma)$ to denote a Gaussian random vector with mean $\mu$ and covariance matrix $\Sigma$, and $U[a,b]$ to denote a random variable drawn uniformly on the interval $[a,b]$. Finally, $P_{v}$, where $v$ is a vector, denotes the orthogonal projection onto $\text{span}(v)$.
3. Stochastic Path Following Compression
We focus on the $L$-layer perceptron [^1] Our algorithm and theory are also applicable to convolutional neural networks, but for simplicity we consider only multi-layer perceptrons (see, e.g.,[30]) $\Phi$ defined above. The aim is to compress the weights of $\Phi$ through quantization or pruning to obtain a network $\widetilde{\Phi}$ with $\Phi(X)\approx\widetilde{\Phi}(X)$. We accomplish this via Algorithm 1 below which we now briefly explain.
Algorithm 1 Stochastic Path Following Compression($\mathcal{T},\theta,C,\Phi,X$)
Parameters: Stochastic operator $\mathcal{T}$, failure threshold $\theta>0$, constant $C\geq 1$.
Input: $L$-layer perceptron $\Phi$ with weight matrices $W^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$, data $X\in\mathbb{R}^{m\times N_{0}}$.
Output: Compressed neural network $\widetilde{\Phi}$.
1:for $i=1$ to $L$ do
2:     Set $X^{(i-1)}=\Phi^{(i-1)}(X)\in\mathbb{R}^{m\times N_{i-1}}$.
3:     Set $\widetilde{X}^{(i-1)}=\widetilde{\Phi}^{(i-1)}(X)\in\mathbb{R}^{m\times N_{i-1}}$.
4:     repeat
5:         Choose a column $w\in\mathbb{R}^{N_{i-1}}$ of $W^{(i)}$.
6:         Run Algorithm 2 to obtain “compressed” weights $q$:
$q=\textsc{Compression}(\mathcal{T},\theta,C,w,X,\widetilde{X}).$
7:     until All columns of $W^{(i)}$ are compressed into $Q^{(i)}$.
8:     Set: $\widetilde{\Phi}^{(i)}(\cdot)=\rho^{(i)}\circ\widetilde{A}^{(i)}\circ\rho^{(i-1)}\circ...\circ\rho^{(1)}\circ\widetilde{A}^{(1)}(\cdot)$, where $\widetilde{A}^{(i)}(x)=Q^{(i)\top}x.$
9:end for
Algorithm 1 uses $m$ data points in $\mathbb{R}^{N_{0}}$ represented by $X\in\mathbb{R}^{m\times N_{0}}$ to guide the compression. The basic idea is to perform the compression one layer at a time, starting with the first. Thus, at layer $i$, we have access to $X^{(i-1)}:=\Phi^{(i-1)}(X)$ and $\widetilde{X}^{(i-1)}:=\widetilde{\Phi}^{(i-1)}(X)$, both in $\mathbb{R}^{m\times N_{i-1}}$, and the goal is to compress $W^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ into $Q^{(i)}\in\mathbb{R}^{N_{i-1}\times N_{i}}$ such that $X^{(i-1)}W^{(i)}\approx\widetilde{X}^{(i-1)}Q^{(i)}$. To accomplish this we use Algorithm 2 to independently, and possibly in parallel, replace each column (neuron) $w\in\mathbb{R}^{N_{i-1}}$ of $W^{(i)}$, entry by entry, with a new neuron $q$ such that $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$. In Algorithm 2, to ensure $X^{(i-1)}w\approx\widetilde{X}^{(i-1)}q$, we proceed sequentially by computing, for $t=1,2,...,N_{i-1}$, the quantized or pruned coefficient $q_{t}$ that aligns the sums $\sum_{j=1}^{t}w_{j}X^{(i-1)}_{j}$ and $\sum_{j=1}^{t}q_{j}\widetilde{X}^{(i-1)}_{j}$. This step forms the core of our method. It balances two competing requirements, namely reducing the error while ensuring that $q_{t}$ lies within the quantization alphabet or has a high likelihood of being set to zero, depending on the application.
Denote the accumulated error at step $t-1$ by $u_{t-1}:=\sum_{j=1}^{t-1}w_{j}X^{(i-1)}_{j}-\sum_{j=1}^{t-1}q_{j}\widetilde{X}^{(i-1)}_{j}$. To balance these requirements, we choose $q_{t}$ such that $\tfrac{1}{C}\cdot u_{t-1}+w_{t}X^{(i-1)}_{t}-q_{t}\widetilde{X}^{(i-1)}_{t}\approx 0$, where $C\geq 1$ controls the trade-off between correcting the prior error and only approximating $w_{t}$. If we were picking $q_{t}$ to minimize $\|\tfrac{1}{C}\cdot u_{t-1}+w_{t}X^{(i-1)}_{t}-q_{t}\widetilde{X}^{(i-1)}_{t}\|$ over $\mathbb{R}$, in the absence of quantization or sparsity constraints, $q_{t}$ is best chosen by projecting $\tfrac{1}{C}\cdot u_{t-1}+w_{t}X^{(i-1)}_{t}$ onto $\widetilde{X}^{(i-1)}_{t}$, yielding $q_{t}=\tfrac{\langle h_{t},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}$, where $h_{t}=Cw_{t}X^{(i-1)}_{t}+u_{t-1}$. However, since our goal is quantization or pruning, we instead set $q_{t}=\mathcal{T}\left(\tfrac{\langle h_{t},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}\right)$, where $\mathcal{T}$ is a stochastic operator that promotes sparsity and/or quantizes its argument. Each neuron $w$ can be handled this way in parallel to compress the entire layer.
In the case of quantization, if $\tfrac{\langle h_{t},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}$ falls outside of the boundaries of the chosen alphabet, the error can become uncontrollable. In [29], where $C=1$, this is prevented by simply adding more elements to the alphabet. In contrast, in order to work with 1-bit alphabets we allow $C\geq 1$, set a threshold $\theta$, and terminate our algorithm if $\left|\tfrac{\langle u_{t-1},\widetilde{X}^{(i-1)}_{t}\rangle}{C\|\widetilde{X}^{(i-1)}_{t}\|^{2}}\right|>\theta$. For instance, if $\{-2K,2K\}$ is the alphabet, then $\theta$ is set to $K$. Meanwhile, for pruning, no such termination is necessary, so $\theta$ is set to infinity. In the remainder of this paper, we show that these modifications allow us to control the error with high probability on the stochasticity of the operator $\mathcal{T}$.
Algorithm 2 Compression($\mathcal{T},\theta,C,w,X,\widetilde{X}$)
Parameters: Stochastic operator $\mathcal{T}$, failure threshold $\theta>0$, constant $C\geq 1$.
Input: Neuron $w=(w_{1},w_{2},\dots,w_{N})^{\top}\in\mathbb{R}^{N}$, data (activations) $X=(X_{1},X_{2},\dots,X_{N})\in\mathbb{R}^{m\times N}$, activations from quantized network $\widetilde{X}=(\widetilde{X}_{1},\widetilde{X}_{2},\dots,\widetilde{X}_{N})\in\mathbb{R}^{m\times N}$.
Output: Compressed weights $q=(q_{1},\dots,q_{N})^{\top}$.
1:Initialize accumulated error $u_{0}=0$.
2:for $t=1$ to $N$ do
3:     if $\left|\tfrac{\langle u_{t-1},\widetilde{X}_{t}\rangle}{C\|\widetilde{X}_{t}\|^{2}}\right|>\theta$ then
4:         Algorithm Fails
5:         return
6:     end if
7:     Compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$
8:     Compute $v_{t}=\tfrac{\langle h_{t},\widetilde{X}_{t}\rangle}{C\|\widetilde{X}_{t}\|^{2}}$
9:     Compute $q_{t}=\mathcal{T}(v_{t})$
10:     Update accumulated error: $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}\widetilde{X}_{t}$
11:end for
4. Theoretical analysis
The primary focus of our theoretical analysis of Algorithm 1 is the error analysis for a single layer of the network. Specifically, we aim to bound the error introduced when compressing the weight matrix of a single layer. Consider the first layer of a neural network, characterized by the weight matrix $W=(W_{1},W_{2},\ldots,W_{N_{1}})\in\mathbb{R}^{N_{0}\times N_{1}}$, where each $W_{i}$ represents a neuron (column) of $W$ [^2] One could also consider any arbitrary layer and work with $X$ being either the corresponding activations of the compressed or original network. Given the input data $X\in\mathbb{R}^{m\times N_{0}}$, Algorithm 1 compresses $W$ into $Q=(Q_{1},Q_{2},\ldots,Q_{N_{1}})$ and the error of this compression is measured in the Frobenius norm,
$\|XW-XQ\|_{F}^{2}=\sum_{i=1}^{N_{1}}\|X(W_{i}-Q_{i})\|_{2}^{2}.$
Since this function is separable with respect to $Q_{1},\ldots,Q_{N_{1}}$, Algorithm 1 processes each column $W_{1},W_{2},\ldots,W_{N_{1}}$ independently to produce $Q_{1},Q_{2},\ldots,Q_{N_{1}}$. To simplify notation, we represent a single column by lowercase letters $w$ and $q$, denoting the uncompressed and compressed versions, respectively. The error analysis for the entire layer thus reduces to bounding $\|X(w-q)\|_{2}^{2}$, which corresponds to the error from Algorithm 2. The remainder of this section is dedicated to proving Theorem 4.2, which provides explicit bounds on the accumulated error at each step $t$. Its proof requires the concept of convex ordering which we will first introduce. Then, we state and prove Theorem 4.2 using a series of technical lemmas, which we also subsequently prove.
4.1. Main Theorem
We now define convex ordering, a notion that will allow us to control random vectors generated by our algorithm by replacing them with Gaussian random vectors.
Definition 4.1.
Let $X,Y$ be $n$-dimensional random vectors such that
$\mathbb{E}f(X)\leq\mathbb{E}f(Y)$
holds for all convex functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, provided the expectations exist. Then $X$ is said to be less than $Y$ in the convex order, denoted by $X\prec_{cx}Y$.
In order to state and prove our main result on the error associated with Algorithm 2 when $X=\widetilde{X}$, we will first introduce an equivalent formulation of our algorithm. Recall that we initialize the accumulated error as $u_{0}=0$ and note that at each step $t$, we first check a failure condition. If the condition is not met, we compute $h_{t}=Cw_{t}X_{t}+u_{t-1}$. Then we use $h_{t}$ to compute $v_{t}=\tfrac{\langle h_{t},X_{t}\rangle}{C\|X_{t}\|^{2}}$ and compute the compressed entry $q_{t}=\mathcal{T}(v_{t})$. Finally, we update the accumulated error $u_{t}=u_{t-1}+w_{t}X_{t}-q_{t}X_{t}$ and move on to the next step. Since the algorithm fails when the failure condition is met at any step, we also have to track whether the algorithm fails at some step. Thus, instead of studying $u_{t}$ directly, we introduce and analyze an the auxiliary quantity $\widetilde{u}_{t}$, which encodes the information about failure. The sequence $\widetilde{u}_{t}$ with $\widetilde{u}_{1}=u_{1}$ is recursively defined via
$\widetilde{u}_{t}$ =$=\begin{cases}0,&\text{if }\left|\frac{\langle\widetilde{u}_{i-1},X_{i}\rangle}{C\|X_{i}\|^{2}}\right|>\theta\text{ for any }i\leq t,\\ \widetilde{u}_{t-1}+w_{t}X_{t}-q_{t}X_{t},&\text{otherwise}.\end{cases}$
[…]
Using convex ordering, we can control the probability $\mathbb{P}\{\text{$|\tfrac{\langle\widetilde{u}_{i-1},X_{i}\rangle}{C\|X_{i}\|^{2}}|\leq\theta$ for all $i$ and }\|\widetilde{u}_{N}\|_{\infty}\leq\kappa\}$ by showing that $\widetilde{u}_{i}$ is dominated by a Gaussian random vector. This result, stated in the theorem below, allows us to control the errors and probabilities for the quantization and pruning applications in Sections 5, 6, and 7. Here we adapt the proof technique from [1].
Theorem 4.2.
Let $C\geq 1$. Assume the weight vector satisfies $\|w\|_{\infty}\leq K$. Further, assume that the operator $\mathcal{T}$ satisfies $\mathbb{E}[\mathcal{T}(v)]=v$ and $|\mathcal{T}(v)-v|\leq M$ for all $v$ with $|v|\leq K+\theta$. Then, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, the resulting $\widetilde{u}_{j}$ satisfies
$\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\beta_{j}I),\quad\text{for }1\leq j\leq N,$
where $\beta_{j}$ is given by
$\beta_{j}=\frac{C\pi M^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2}.$
Proof.
First by Lemma 4.4, we deduce $\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\Sigma_{j})$, where $\Sigma_{j}$ is defined as in Lemma 4.4. Further by Lemma 4.5, we have $\Sigma_{j}\preceq\beta_{j}I$. We then apply Lemmas 4.3, c and a to complete the proof. ∎
4.2. Technical Lemmas
In this subsection, we present the lemmas used for the proof of Theorem 4.2. We start with a lemma collecting various results on convex ordering, with the proofs in [1, 22].
Lemma 4.3.
The following all hold:
  1. a.
    If $X\prec_{cx}Y$ and $Y\prec_{cx}Z$, then $X\prec_{cx}Z$.
  2. b.
    If $X\prec_{cx}Y$, then for any linear transformation $M$ on $\mathbb{R}^{n}$, we have $MX\prec_{cx}MY$.
  3. c.
    If $A$ and $B$ are two positive semi-definite matrices and $A\preceq B$, then $\mathcal{N}(0,A)\prec_{cx}\mathcal{N}(0,B)$.
  4. d.
    Let $X$, $Y$, $W$, and $Z$ be random vectors and suppose that $X$ and $Y$ live on the same probability space. Let $W$ and $Z$ be independent and suppose that $X\prec_{\mathrm{cx}}W$ and $(Y-X)|X\prec_{\mathrm{cx}}Z$. Then $Y\prec_{\mathrm{cx}}W+Z$.
  5. e.
    Let $X$ be a real-valued random variable with $\mathbb{E}X=0$ and $|X|\leq C$. Then $X\prec_{\mathrm{cx}}\mathcal{N}\bigl{(}0,\frac{\pi C^{2}}{2}\bigr{)}$.
  6. f.
    Let $X$ be a symmetric random vector in $\mathbb{R}^{n}$, i.e., $X$ and $-X$ have the same distribution, and let $A\subset\mathbb{R}^{n}$ be a symmetric set (i.e., $x\in A\Rightarrow-x\in A$). Define $X^{\prime}$ by
    $X^{\prime}=\begin{cases}X,&\text{if }X\in A^{c},\\ 0,&\text{otherwise.}\end{cases}$
    Then $X^{\prime}\prec_{cx}X$.
  7. g.
    Let $X$ be an $n$-dimensional random vector such that $X\prec_{\mathrm{cx}}\mathcal{N}(\mu,\sigma^{2}I)$, and let $\alpha>0$. Then
    $\mathbb{P}\bigl{(}\|X-\mu\|_{\infty}\leq\alpha\bigr{)}\geq 1-\sqrt{2}ne^{-\frac{\alpha^{2}}{4\sigma^{2}}}.$
    In particular, if $\alpha=2\sigma\sqrt{\log(\tfrac{\sqrt{2}n}{\gamma})}$ with $\gamma\in(0,1]$, we have
    $\mathbb{P}\big{(}\|X-\mu\|_{\infty}\leq 2\sigma\sqrt{\log(\tfrac{\sqrt{2}n}{\gamma})}\big{)}\geq 1-\gamma.$
Using Lemma 4.3 allows to deduce the following result on the distribution of $\widetilde{u}_{j}$. We use a proof technique similar to that in [1].
Lemma 4.4.
Under the same assumptions as in Theorem 4.2, by running Algorithm 2 with $w\in\mathbb{R}^{N_{0}}$ and $X\in\mathbb{R}^{m\times N_{0}}$, we obtain $\widetilde{u}_{j}$ such that:
$\widetilde{u}_{j}\prec_{cx}\mathcal{N}(0,\Sigma_{j})\text{ for $1\leq j\leq N$},$
where $\Sigma_{0}=0$ and $\Sigma_{j},j\geq 1$ is defined recursively as
$\Sigma_{j}=\Big{(}I-\dfrac{P_{X_{j}}}{C}\Big{)}\Sigma_{j-1}\Big{(}I-\dfrac{P_{X_{j}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{j}X_{j}^{\top}.$
Proof.
We use induction. For $j=0$, the statement is trivial. Assume the statement is true for $j=t-1$. Define $u^{\prime}_{t-1}$ as
$u^{\prime}_{t-1}=\left\{\begin{split}&\widetilde{u}_{t-1},&\text{ if $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$,}\\ &0,&\text{ otherwise.}\end{split}\right.\quad$
The random variable $\widetilde{u}_{t-1}$ is symmetric, so $u^{\prime}_{t-1}\prec_{cx}\widetilde{u}_{t-1}$ by Lemma 4.3, f. If $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$, then the algorithm does not fail and we have
$\widetilde{u}_{t}$ $=\widetilde{u}_{t-1}+w_{t}X_{t}-q_{t}X_{t}$
$=\widetilde{u}_{t-1}+w_{t}X_{t}-\widetilde{v}_{t}X_{t}+(\widetilde{v}_{t}-q_{t})X_{t}$
$=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\widetilde{u}_{t-1}+(\widetilde{v}_{t}-\mathcal{T}(\widetilde{v}_{t}))X_{t}$
due to the definition of $\widetilde{v}_{t}$ and $q_{t}$. If $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|>\theta$, then we have $\widetilde{u}_{t}=0$ by definition. Thus we can express $\widetilde{u}_{t}$ using $u^{\prime}_{t-1}$ as
$\widetilde{u}_{t}=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u^{\prime}_{t-1}+R_{t}X_{t}$
where
$R_{t}=\left\{\begin{split}&\widetilde{v}_{t}-\mathcal{T}(\widetilde{v}_{t}),&\text{ if $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$,}\\ &0,&\text{ else.}\end{split}\right.\quad$
When $|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$, we have
$|\widetilde{v}_{t}|=\Big{|}\dfrac{\langle\widetilde{h}_{t},X_{t}\rangle}{C\|X_{t}\|^{2}}\Big{|}=\Big{|}w_{t}+\dfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}\Big{|}\leq K+\theta.$
By the assumption about $\mathcal{T}$, $R_{t}$ is mean zero and $|R_{t}|\leq M$. We then use e from Lemma 4.3, to deduce $R_{t}\prec_{cx}\mathcal{N}(0,\tfrac{\pi M^{2}}{2})$. Further applying Lemma 4.3, b, we obtain
$R_{t}X_{t}\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
Since $u^{\prime}_{t-1}\prec_{cx}\widetilde{u}_{t-1}\prec_{cx}\mathcal{N}(0,\Sigma_{t-1})$, we have
$\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}u^{\prime}_{t-1}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Big{)}.$
Finally, by Lemma 4.3, d, we conclude
$\widetilde{u}_{t}\prec_{cx}\mathcal{N}\Big{(}0,\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}+\dfrac{\pi M^{2}}{2}X_{t}X_{t}^{\top}\Big{)}.$
This completes the induction. ∎
Having controlled our object of interest by a Gaussian random vector with a known covariance matrix, $\Sigma_{j}$ defined in Lemma 4.4, we next control the covariance matrix itself.
Lemma 4.5.
Let $C\geq 1$, then for the above defined $\Sigma_{j}$, we have
$\Sigma_{j}\preceq\beta_{j}I,$
where $\beta_{0}=0$ and $\beta_{j}=\dfrac{C\pi M^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2},$ when $j\geq 1$.
Proof.
The base case $j=0$ is obvious. Let us now proceed by induction. Suppose $\Sigma_{t-1}\preceq\beta_{t-1}I$. Then
$\Sigma_{t}$ $=\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}\Sigma_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}+\dfrac{\pi M^{2}X_{t}X_{t}^{\top}}{2}$
$\preceq\beta_{t-1}\Big{(}I-\dfrac{P_{X_{t}}}{C}\Big{)}^{2}+\dfrac{\pi M^{2}X_{t}X_{t}^{\top}}{2}$
$=\beta_{t-1}\Big{(}I+\dfrac{1}{C^{2}}\dfrac{X_{t}X_{t}^{\top}}{\|X_{t}\|^{2}}-2\dfrac{X_{t}X_{t}^{\top}}{C\|X_{t}\|^{2}}\Big{)}+\dfrac{\pi M^{2}X_{t}X_{t}^{\top}}{2}$
$\preceq\beta_{t}\Big{(}I+\dfrac{1}{C^{2}}\dfrac{X_{t}X_{t}^{\top}}{\|X_{t}\|^{2}}-2\dfrac{X_{t}X_{t}^{\top}}{C\|X_{t}\|^{2}}\Big{)}+\dfrac{\pi M^{2}X_{t}X_{t}^{\top}}{2}$
$=\beta_{t}I+\dfrac{1}{\|X_{t}\|^{2}}\Big{(}\beta_{t}\Big{(}\dfrac{1}{C^{2}}-\dfrac{2}{C}\Big{)}+\dfrac{\pi M^{2}}{2}\|X_{t}\|^{2}\Big{)}X_{t}X_{t}^{\top}$
$=\beta_{t}I+\dfrac{\pi M^{2}}{\|X_{t}\|^{2}}\Big{(}\dfrac{C}{2}\max\limits_{1\leq i\leq t}\|X_{i}\|^{2}\Big{(}\dfrac{1-2C}{C^{2}}\Big{)}+\dfrac{1}{2}\|X_{t}\|^{2}\Big{)}X_{t}X_{t}^{\top}.$
Since $C\geq 1$, we know $1-2C\leq 0$, then we can further deduce
$\Sigma_{t}$ $\preceq\beta_{t}I+\dfrac{\pi M^{2}}{\|X_{t}\|^{2}}\Big{(}\dfrac{C}{2}\|X_{t}\|^{2}\Big{(}\dfrac{1-2C}{C^{2}}\Big{)}+\dfrac{1}{2}\|X_{t}\|^{2}\Big{)}X_{t}X_{t}^{\top}$
$=\beta_{t}I+\Big{(}\dfrac{\pi M^{2}}{2}\Big{(}\dfrac{1}{C}-2\Big{)}+\dfrac{\pi M^{2}}{2}\Big{)}X_{t}X_{t}^{\top}$
$\preceq\beta_{t}I+\Big{(}-\dfrac{\pi M^{2}}{2}+\dfrac{\pi M^{2}}{2}\Big{)}X_{t}X_{t}^{\top}$
$=\beta_{t}I.$
This completes the induction. ∎
5. One-bit Quantization
We are now ready to first describe the selection of $\mathcal{T}$ and $\theta$ for one-bit quantization and then provide the corresponding error analysis for running Algorithm 1 on a single layer of a neural network. Notably, our analysis extends with minimal additional effort to finite multi-bit quantization, a consideration we leave to the interested reader.
5.1. Choice of $\mathcal{T}$ and $\theta$ for one-bit quantization
Suppose we want to quantize an $L$-layer perceptron which has all its weights bounded in absolute value by $K>0$. We use $\mathcal{A}=\{-2K,2K\}$ as the alphabet. Then, we can choose $\mathcal{T}$ to be the stochastic scalar quantizer $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$, with
$\mathcal{Q}(z)=\begin{cases}2K&\text{with probability $p_{z}$ if $|z|\leq 2K$}\\ -2K&\text{with probability $1-p_{z}$ if $|z|\leq 2K$}\\ \text{sgn}(z)\cdot 2K&\text{if $|z|>2K$},\end{cases}$
where $p_{z}=\tfrac{1}{2}+\tfrac{z}{4K}$. Moreover, we choose the threshold $\theta=K$. It is straightforward to verify that the operator $\mathcal{Q}$ satisfies the assumptions in Theorem 4.2 with $M=4K$.
5.2. Error analysis
Consider one layer of a neural-network, and denote its action by $\Phi(x)=\rho(W^{\top}x)$, where $x\in\mathbb{R}^{N_{0}}$ represents the input data or activations, $W\in\mathbb{R}^{N_{0}\times N_{1}}$ is the weight matrix and $\rho$ is the ReLU activation function, or any (say) 1-Lipschitz function. Now, we use Algorithm 1 to quantize $\Phi$ using $X\in\mathbb{R}^{m\times N_{0}}$, where each row represents a data point. The following holds.
Proposition 1.
Let $C\geq 1$. Suppose that $Q$ is the quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic quantizer $\mathcal{Q}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$
with probability greater than
$1-N_{1}\sum_{t=1}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}-\sqrt{2}mN_{1}N_{0}^{-p},$
where $X_{t}$ represents the $t$-th column of $X$.
Proof.
Let $\kappa:=4K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$. Consider any neuron (column) $w\in\mathbb{R}^{\mathbb{N}_{0}}$ from $W$. Let $q$ and $\widetilde{u}_{i}$ be as defined in Subsection 4.1. We run Algorithm 2 independently for each $w$. We can deduce
[…]
$=\mathbb{P}\big{(}\text{$|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|\leq\theta$ for all $t$ and }\|\widetilde{u}_{N}\|_{\infty}\leq\kappa\big{)}$
$=1-\mathbb{P}(\text{$|\tfrac{\langle\widetilde{u}_{t-1},X_{t}\rangle}{C\|X_{t}\|^{2}}|>K$ for some $t$ or }\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)$
$\geq 1-\sum_{t}\mathbb{P}\Big{(}\Big{|}\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\Big{|}>K\Big{)}-\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa).$
By Theorem 4.2 and Lemma 4.3, b, we have
$\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\prec_{cx}\mathcal{N}\Big{(}0,\dfrac{\beta_{t}}{C^{2}\|X_{t+1}\|^{2}}\Big{)},$
where $\beta_{t}=8C\pi K^{2}\max_{1\leq i\leq t}\|X_{i}\|^{2}.$ Applying Lemma 4.3, g, we obtain
$\mathbb{P}\Big{(}\Big{|}\dfrac{\langle\widetilde{u}_{t},X_{t+1}\rangle}{C\|X_{t+1}\|^{2}}\Big{|}>K\Big{)}\leq\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t+1}\|^{2}}{32\pi\max_{1\leq i\leq t}\|X_{i}\|^{2}}}\Big{)}.$ (1)
For $\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)$, we use the second part of Lemma 4.3, g with $\gamma=\sqrt{2}mN_{0}^{-p}$ to get
$\mathbb{P}(\|\widetilde{u}_{N_{0}}\|_{\infty}>\kappa)\leq\sqrt{2}mN_{0}^{-p}.$ (2)
Thus using (1) and (2), we obtain
[…]
$\geq 1-\sum_{t=2}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}-\sqrt{2}mN_{0}^{-p}.$
This implies that
$\mathbb{P}(\|\rho(Xw)-\rho(Xq)\|_{\infty}\leq\kappa)$ (3)
$\geq 1-\sum_{t=2}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{32\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}-\sqrt{2}mN_{0}^{-p},$
by the definition of $\widetilde{u}_{N_{0}}$ and the the Lipschitz property of the activation function $\rho$. Inequality (3) holds true for each column of $W\in\mathbb{R}^{m\times N_{1}}$. Taking a union bound on the probability completes the proof. ∎
Remark 1.
This result demonstrates that by choosing $C$ proportional to $\log(N_{0}N_{1})$ the error associated with quantization scales as ${\operatorname{polylog}N_{0}}$, with high probability. This stands in sharp contrast to the error from naive rounding algorithms, where each entry of $w$ is rounded to the nearest element of $\mathcal{A}$. In the naive case, the error scales as $\sqrt{N_{0}}$.
Remark 2.
For multi-layer analysis, we can use the ‘data alignment’ technique introduced in [29], which would allow establishing an error bound for the entire neural network. We leave this to the reader.
6. Pruning
In this section, we specify the choice of $\mathcal{T}$ and the threshold $\theta$ used for pruning, and we present the error analysis of applying Algorithm 1 to a single layer of a neural network.
6.1. Choice of $\mathcal{T}$ and $\theta$ for Pruning
Suppose our layer weights are bounded by $K>0$ in absolute value. Given $c>0$, our goal is to encourage weights below $cK$ to be set to 00 while keeping weights above $cK$ unchanged. To achieve this, we set $\mathcal{T}$ to the stochastic operator $\mathcal{S}$, with
$\mathcal{S}(z)=\begin{cases}z&\text{ if $|z|>cK$,}\\ \xi_{z}&\text{ if $|z|\leq cK$.}\end{cases}$ (4)
Here $\xi_{z}$ is a random variable that is set to 0 with probability $1-\tfrac{1}{(c+1/2)K}|z|$ and is sampled from $\operatorname{sgn}(z)\cdot{U}[cK,K]$ with probability $\tfrac{1}{(c+1/2)K}|z|$. Additionally, we set $\theta$ to be $\infty$. It is straightforward to verify that the operator $\mathcal{S}$ satisfies the assumptions in Theorem 4.2 with $M$ being $K$.
6.2. Error analysis
Consider again $\Phi(x)=\rho(W^{\top}x)$, where $x\in\mathbb{R}^{N_{0}}$ is the input data or activation, $W\in\mathbb{R}^{N_{0}\times N_{1}}$ is the weight matrix and $\rho$ is the ReLU (or any 1-Lipschitz) activation function. We apply Algorithm 1 to prune $\Phi$ using $X\in\mathbb{R}^{m\times N_{0}}$, where each row represents a data point. We can then deduce the following result.
Proposition 2.
Let $C\geq 1$. Suppose that $Q$ is the pruned weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic pruning operator $\mathcal{S}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq K\sqrt{2C\pi p\log N_{0}}\max_{1\leq i\leq N_{0}}\|X_{i}\|$
with probability greater than $1-\sqrt{2}mN_{1}N_{0}^{-p}$, where $X_{t}$ represents the $t$-th column of $X$.
Proof.
Consider any neuron (column) $w\in\mathbb{R}^{\mathbb{N}_{0}}$ from $W$. Let $q$ and $\widetilde{u}_{i}$ be defined as in Subsection 4.1. We run Algorithm 2 independently for each $w$. In this case, $\widetilde{u}_{i}=u_{i}$ for all $i$.
It is straightforward to verify that the assumptions in Theorem 4.4 are satisfied with $M=K$. Thus, from Theorem 4.2, we deduce the accumulated error satisfies
$u_{j}\prec_{cx}\mathcal{N}(0,\beta_{j}I)\text{ for $1\leq j\leq N_{0}$},$
where $\beta_{j}=\tfrac{C\pi K^{2}}{2}\max_{1\leq i\leq j}\|X_{i}\|^{2}.$
Applying Lemma 4.3, g with $\gamma=\sqrt{2}mN_{0}^{-p}$, we obtain
$\mathbb{P}(\|u_{N_{0}}\|_{\infty}>K\sqrt{2C\pi p\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|)\leq\sqrt{2}mN_{0}^{-p}.$
This implies
$\mathbb{P}(\|\rho(Xw)-\rho(Xq)\|_{\infty}>K\sqrt{2C\pi p\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|)\leq\sqrt{2}mN_{0}^{-p}.$ (5)
The inequality (5) holds for any column $w$ of $W\in\mathbb{R}^{N_{0}\times N_{1}}$. Applying a union bound on the probability completes the proof. ∎
7. Quantization with pruning
In this section, we show that our methods can handle combinations of operators, and provide the choices for $\mathcal{T}$ and the threshold $\theta$ used in quantization with pruning, followed by an error analysis for applying Algorithm 1 to a single layer of a neural network.
7.1. Choice of $\mathcal{T}$ and $\theta$ for Quantization with Pruning
Consider an $L$-layer perceptron with all weights bounded by $K>0$ in absolute value. Given $c>0$, and alphabet $\mathcal{A}=\{-2K,0,2K\}$, our goal is to encourage weights below $cK$ to be set to 00 and then quantize the weights using $\mathcal{A}$. Let $\mathcal{S}$ be the operator defined in Section 6. Also, we define the following stochastic rounding operator $\mathcal{Q}:\mathbb{R}\rightarrow\mathcal{A}$
$\mathcal{Q}(z)=\begin{cases}\lfloor\tfrac{z}{2K}\rfloor\cdot 2K&\text{with probability $p_{z}$ if $|z|\leq 2K$}\\ \left(\lfloor\tfrac{z}{2K}\rfloor+1\right)\cdot 2K&\text{with probability $1-p_{z}$ if $|z|\leq 2K$}\\ \text{sgn}(z)\cdot 2K&\text{if $z>2K$},\end{cases}$
where $p_{z}=1-\tfrac{z}{2K}+\lfloor\tfrac{z}{2K}\rfloor$. We then define $\mathcal{T}=\mathcal{Q}\circ\mathcal{S}$ with the threshold $\theta=K$. As before, it is straightforward to verify that the operator $\mathcal{S}$ satisfies the assumptions of Theorem 4.2 with $M=2K$.
7.2. Error analysis
With the same notation as before, we can deduce the following result.
Proposition 3.
Let $C\geq 1$. Suppose that $Q$ is the pruned and quantized weight matrix resulting from Algorithm 1, applied to $\Phi$ with the stochastic operator $\mathcal{Q}\circ\mathcal{S}$. Then, the resulting error satisfies
$\max_{i,j}|\rho(XW)_{ij}-\rho(XQ)_{ij}|\leq 2K\sqrt{2\pi Cp\log N_{0}}\max_{1\leq i\leq N}\|X_{i}\|$
with probability greater than
$1-N_{1}\sum_{t=1}^{N_{0}}\sqrt{2}\exp\Big{(}{-\dfrac{C\|X_{t}\|^{2}}{8\pi\max_{1\leq i\leq t-1}\|X_{i}\|^{2}}}\Big{)}-\sqrt{2}mN_{1}N_{0}^{-p},$
where $X_{t}$ represents the $t$-th column of $X$.
The proof of this Proposition is essentially identical to that of Proposition 1, so we omit the details.
Acknowledgment
We gratefully acknowledge partial support by National Science Foundation Grants DMS-2012546 and DMS-2410717.